Foam fingers sold for $5. Spirit bracelets for $4. You sold 40 more foam fingers than bracelets and earned $965. Write a system of equations to describe this situation.
Solve this system using as least 2 different methods. Explain each method.

Answer by HyperBrain(694)

Let x=no. of foam fingers sold
y=no. of spirit bracelets sold
then
5x=amount earned on foam fingers
4y=amount sold on spirit bracelets
You sold 40 more foam fingers than bracelets;: so,
x=y+40
you earned $965. so,
5x+4y=965
substituting x=y+40,
5(y+40)+4y=965
5y+200+4y=965
9y+200=965
9y=765
y=85
x=y+40=85+40=125
Therefore, you sold 125 foam fingers and 85 spirit bracelets.
//You can contact me at adrianvidal07@hotmail.com Don't spam.
Power up,
HyperBrain!
Stop this mess. Visit:
http://uselessplastic.wordpress.com
http://uselessplastic.multiply.com